# GaMeS: Mesh-Based Adapting and Modification of Gaussian Splatting

### v3

@@ 3 GaMeS : Mesh-Based Gaussian Splatting
GaMeS (our - pseudo-mesh) 34.74 25.94 35.88 36.58 30.50 30.66 37.84 35.41
RIP-NeRF 34.84 24.89 33.41 34.19 28.31 30.65 35.96 32.23
Table 1: Quantitative comparisons (PSNR) on a NeRF-Synthetic dataset showing that GaMeS gives comparable results with other models.
Initial Position GaMeS Only Gaussians means on mesh Refer to caption
Figure 9: Compared to a method solely based on Gaussian averaging dependent on position, we observe that, following rotation and stretching the surface, Gaussians should dynamically adjust their scale and rotation to seamlessly conform to the change as in the case of GaMeS model.
@@
GaMeS parameterize all Gaussian components by verities of mesh in such a way that they lie on meshes. Therefore, when we modify mesh positions, we automatically adapt the means and covariances of relevant Gaussians. Such an approach allows us to work with transformation. To better visualize such properties, we compare our model with a solution where the Gaussian components’ centers are placed on mesh faces but do not parameterize the covariance matrix. In Fig. 8 we can see that GaMeS Gaussian component adapts automatically to mesh modification.
@@ 4 Experiments
Here, we delineate implementation details, describe the used datasets, and the reasons for their selection. Then, we describe three scenarios: (1) when we have an existing mesh, (2) when we use pseudo-mesh estimated during training, and (3) when we use FLAME initialization. In the end, we show inference time both on static and animated scenes.
4.1 Implementation details and data description
                              PSNR $\uparrow$
@@ 4.1 Implementation details and data description
R-SuGaR-15K 23.14 - 25.36 24.70 - 30.03 27.62 29.56 30.51
GaMeS (Our-pseudo-mesh) 24.99 21.27 27.22 26.54 22.39 31.52 28.92 31.12 32.09
@@
Table 2: The quantitative comparisons of reconstruction capability (PSNR) on Mip-NeRF360 dataset. R-SuGaR-15K uses 1M vertices. GaMeS uses pseudo-mesh.
The time taken by the optimization process for our model varied from a few up to no more than sixty minutes, as it depends on factors like scene complexity, dataset characteristics, the number of Gaussians per face, and vertex fine-tuning decisions. Nonetheless, all experiments were conducted within a reasonable time frame, underscoring the efficiency of our approach across diverse scenarios. The supplementary material provides all the relevant details, and the source code is available in Github [^1] https://github.com/waczjoan/gaussian-mesh-splatting.
We illustrated the fundamental gains of GaMeS model through experiments with three distinct datasets.
@@
Synthetic Dataset: A NeRF-provided eight geometrically complex objects with realistic non-Lambertian materials (Mildenhall et al., 2020). In addition to containing images of the objects, the dataset also incorporates corresponding meshes. We leveraged this dataset for our initial experiments to unequivocally showcase the capabilities of our model and underscore its flexibility in mesh editing.
@@
Mip-NeRF360 Dataset: A dataset comprising five outdoor and four indoor scenes, each featuring intricate central objects or areas against detailed backgrounds (Barron et al., 2022). We used this dataset to demonstrate the adaptability of our model in handling dense scenes, employing a non-conventional strategy during the initial training stage in mesh preparation.
@@
Human Faces Dataset: A modified subset of the Stirling/ESRC 3D Face Database [^2] https://pics.stir.ac.uk/ESRC/index.htm, that includes images of six people, generated using Blender software from various perspectives, consistently excluding the backs of heads. Only one expression is assigned for each face. Notably, the dataset lacks corresponding mesh representations for the depicted person. This dataset is used to compare GaMeS with the NeRFlame model (Zając et al., 2023), mainly within the human-mesh fitting task.
4.2 Scenario I: Model with existing mesh
@@
In this scenario, we utilize the provided mesh, incorporating Gaussians by strategically placing them onto its surface. We also allowed for moving vertices. Our initial experiments used the Synthetic Dataset (Mildenhall et al., 2020), incorporating shared meshes. Table 1 presents PSNR metrics demonstrating the competitiveness of GaMeS approach with existing methods. Additional numerical comparisons are available in the supplementary material. Notably, a significant contribution lies in the ease with which renders can be modified manually or in an automated process. Fig. 1 shows examples of reconstruction and simple object animations.
Figure 10: The main idea of fitting human mesh and how it can be transformed into a different person.
4.3 Scenario II: A model without mesh
When we do not have mesh, we train GS with flat Gaussians. Then, we construct the pseudo-mesh. Table 1 shows a comparison with other models on the Synthetic Dataset. Our approach is also capable of modeling unbounded scenes. In order to achieve that, we first create an initial pseudo-mesh and then adjust it during the training. Results in Table 2 demonstrate that such modeling is feasible and reasonably comparable to other existing models. Particularly, the results show parity with the GS-30K and R-SuGaR-15K models.
As anticipated, we observe scene reconstruction, including both objects and details, as can be seen in Fig. 2. Nevertheless, backgrounds and elements with large mesh faces may exhibit a lower-quality appearance. However, this does not hinder the feasibility of GaMeS model for reconstruction and modification.
4.4 Scenario III: GaMeS with FLAME as Init Mesh
In this experiment, we show the possibility of fitting mesh using an Init Mesh, which was not acquired directly from our data in contrast to previous scenarios. Here, we relied on the Faces Dataset as it provides such required experimental data. To initialize the meshes, we used the FLAME (Li et al., 2017) framework, generating fully controllable meshes of human faces. Consequently, we could highlight the key advantage of GaMeS model, i.e., the ability to modify objects. We used the official implementation of the FLAME, with the number of parameters suggested in RingNet (Sanyal et al., 2019) In addition, we implemented an extra detail learning parameter to better represent hair or T-shirts.
In Fig. 10, we show the main idea behind our approach. Initially, a human mesh is generated using the FLAME model. Thanks to accurately adjusting the mesh, we can apply it, for example, to another person. By utilizing the FLAME parameters, we can modify various aspects, especially facial expressions, to generate facial expressions like smiling, sadness, eye closure, mouth opening, and flirty or disgusted expressions. The possibilities of reconstruction and modifications for all six faces are depicted in Fig. 11. All expressions were chosen randomly as our main goal was to show GaMeS model’s ability to render objects as well as its modifications/animations.
PSNR $\uparrow$
@@ 4.4 Scenario III: GaMeS with FLAME as Init Mesh
The comparison of the mesh fitting capability between the GaMeS with the NeRFlame model is shown in Fig. 12. Notably, the faces are distinctly delineated and closely resemble the modeled objects, capturing not only postures but also expressions and intricate details. Table 3 shows the results of using PSNR metrics on six faces. Comparisons are made between the two baseline NeRF models, GS, and the models created from them, i.e., the corresponding NeRFlame and GaMeS with animation capabilities.
Figure 12: During the training of GaMeS, we simultaneously model FLAME mesh. Here, we present the meshes fitted by GaMeS compared to fitted mesh by NeRFlame model.
@@ 5 Conclusion
This paper presents a new GS-based model called GaMeS that leverages two main concepts, i.e., classical mesh and vanilla GS. Here, we represent a 3D scene with a set of Gaussian components lying on the mesh. We can use the initial mesh as an input or estimate it during training. Moreover, we parameterize the Gaussian components by the mesh vertices, allowing single-stage training. In inference, we obtain GS, which can be modified in real-time by manipulating mesh components. As a result, when the mesh changes, all Gaussian components automatically adapt to a new shape.
@@
Limitations GaMeS allows real-time modifications, but artifacts appear in the case of significant changes in the case of meshes with large faces. In practice, large faces should be divided into smaller ones. How to change Gaussian components in GaMeS when mesch faces are split is not apparent.
Societal impact Our GaMeS contributes to lowering the computational complexity of 3D modeling, therefore reducing the energy consumption of such models.
Appendix A Appendix
The appendix to the paper contains excerpts related to NeRF Synthetic and a continuation of the numerical results expounded in the main document.
PSNR $\uparrow$
Number of Gaussians Chair Drums Lego Mic Materials Ship Hotdog Ficus
1 29.32 25.18 32.74 32.48 25.37 27.45 32.10 Z32.05
5 30.70 25.35 33.21 32.49 25.29 29.29 33.40 32.29
10 31.11 25.36 OOM 32.29 25.25 29.56 33.65 32.35
pseudo-mesh 35.34 25.93 35.42 35.34 30.01 29.58 37.65 34.81
SSIM $\uparrow$
Chair Drums Lego Mic Materials Ship Hotdog Ficus
1 0.934 0.940 0.972 0.984 0.916 0.879 0.959 0.978
5 0.950 0.942 0.975 0.984 0.914 0.894 0.968 0.978
10 0.951 0.942 OOM 0.983 0.913 0.896 0.970 0.979
pseudo-mesh 0.986 0.952 0.982 0.991 0.958 0.886 0.985 0.987
LPIPS $\downarrow$
Chair Drums Lego Mic Materials Ship Hotdog Ficus
1 0.066 0.052 0.027 0.012 0.060 0.133 0.062 0.018
5 0.050 0.049 0.022 0.012 0.061 0.107 0.042 0.018
10 0.044 0.049 OOM 0.012 0.063 0.102 0.038 0.019
pseudo-mesh 0.012 0.039 0.015 0.006 0.033 0.105 0.019 0.012
Table 4: The quantitative comparisons (PSNR / SSIM / LPIPS) on Synthetic dataset with original meshes or pseudo-mesh, with white background. OOM - CUDA out of memory. In this experiment we used Tesla V100-SXM2-32GB GPU.
A.1 Number of Gaussians per face
Here, we show the results of the experiments for the NeRF Synthetic dataset with a white background and original mesh. The dependence of the PSNR, SSIM and LPIPS result on the number of Gaussians per face is shown in Table 4.
Mostly, as a consequence of increasing the number of Gaussians, we get better results, and we can observe that a single Gaussian per face is insufficient. Materials and Lego has a very dense mesh, consequently positioning an excessive number of Gaussians within a confined space does not yield substantial improvements in the results.
Initial Mesh Subdivided Mesh
$k=1$ $k=10$ $k=5$
PSNR = 29.32 PSNR = 31.11 PSNR = 33.01
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 13: A qualitative assessment delineating the impact of varying Gaussian quantities per face, discerned between the initial mesh and its subdivided counterpart.
Initial Mesh Subdivided Mesh k=1 k=10 k=5
Figure 14: A juxtaposition of outcomes is conducted for a model varying in Gaussian counts per face and with a mesh subdivided. GaMeS with good mesh initialization is able to capture fine details.
However, experiments have shown that choosing a fixed number of Gaussians per face mesh is inefficient when the mesh contains different face sizes – in particular, it is difficult to cover large faces. Therefore, we decided to split the large facets. Splitting large facets, even using fewer Gaussians per face, allows for significantly better results and detail capturing. The effects are shown in Fig. 13 (initial positions) and Fig. 14 (modifications).
A.2 Modification
Illustrating the reconstruction and modification through the application of the GaMeS model is shown in Fig. 15. GaMeS allows the creation of lifelike alterations, such as elevating the excavator blade or imaginative and surrealistic bends.
In fact, modification depends only on user skills. Fig. 16 shows various modifications of the ficus such as: a) scaling b) modifying a certain part – like bending branches, or modifying the whole object – dancing ficus. (c) ignoring a certain part. All images are generated from the view of the same camera. Note that Gaussians represent color from two sides. Therefore, after removing the soil from the pot, we can still see the inner side of the pot.
Fig. 17 shows that modifications are possible even using only pseudo-mesh.
Before modification   After modification
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 15: An example of reconstruction and modification using the GaMeS model. The model enables more realistic modifications like lifting the excavator blade as well as surrealistic drums bending.
Figure 16: Demonstrations of potential modifications facilitated by games include: a) scaling, b) transforming either specific components or entire objects, and c) selectively disregarding specific sections.
Figure 17: An example of modification using the GaMeS model with pseudo-mesh.
A.3 Extension of numerical results from main paper
Here, we present extensions of experiments proposed in the main part.
PSNR $\uparrow$
Outdoor scenes Indoor scenes
bicycle flowers garden stump treehill room counter kitchen bonsai
Static
Plenoxels 21.91 20.10 23.49 20.66 22.25 27.59 23.62 23.42 24.66
INGP-Big 22.17 20.65 25.07 23.47 22.37 29.69 26.69 29.48 30.69
Mip-NeRF360 24.37 21.73 26.98 26.40 22.87 31.63 29.55 32.23 33.46
GS - 7K 23.60 20.52 26.25 25.71 22.09 28.14 26.71 28.55 28.85
GS - 30K 25.25 21.52 27.41 26.55 22.49 30.63 28.70 30.32 31.98
Editable
R-SuGaR-15K 22.91 - 25.29 24.55 - 29.95 27.47 29.38 30.42
GaMeS (Our-pseudo-mesh) 24.99 21.27 27.22 26.54 22.39 31.52 28.92 31.12 32.09
SSIM $\uparrow$
Outdoor scenes Indoor scenes
bicycle flowers garden stump treehill room counter kitchen bonsai
Static
Plenoxels 0.496 0.431 0.606 0.523 0.509 0.842 0.759 0.648 0.814
INGP-Big 0.512 0.486 0.701 0.594 0.542 0.871 0.817 0.858 0.906
Mip-NeRF360 0.685 0.583 0.813 0.744 0.632 0.913 0.894 0.920 0.941
GS - 7K 0.675 0.525 0.836 0.728 0.598 0.884 0.873 0.900 0.910
GS - 30K 0.771 0.605 0.868 0.775 0.638 0.914 0.905 0.922 0.938
Editable
R-SuGaR-15K 0.631 - 0.771 0.681 - 0.909 0.890 0.907 0.933
GaMeS (Our-pseudo-mesh) 0.757 0.596 0.860 0.766 0.628 0.916 0.904 0.923 0.939
LPIPS $\downarrow$
Outdoor scenes Indoor scenes
bicycle flowers garden stump treehill room counter kitchen bonsai
Static
Plenoxels 0.506 0.521 0.386 0.503 0.540 0.419 0.441 0.447 0.398
INGP-Big 0.446 0.441 0.257 0.421 0.450 0.261 0.306 0.195 0.205
Mip-NeRF360 0.301 0.344 0.170 0.261 0.339 0.211 0.204 0.127 0.176
GS - 7K 0.318 0.417 0.153 0.287 0.404 0.272 0.254 0.161 0.244
GS - 30K 0.205 0.336 0.103 0.210 0.317 0.220 0.204 0.129 0.205
Editable
R-SuGaR-15K 0.349 - 0.218 0.336 - 0.243 0.234 0.166 0.219
GaMeS (Our-pseudo-mesh) 0.216 0.342 0.111 0.219 0.332 0.221 0.203 0.129 0.205
Table 5: The quantitative comparisons (PSNR / SSIM / LPIPS) on the Mip-NeRF360 dataset. R-SuGaR-15K, with the number of 200K vertices.
PSNR $\uparrow$
Chair Drums Lego Mic Materials Ship Hotdog Ficus
Static
NeRF 33.00 25.01 32.54 32.91 29.62 28.65 36.18 30.13
VolSDF 30.57 20.43 29.46 30.53 29.13 25.51 35.11 22.91
Ref-NeRF 33.98 25.43 35.10 33.65 27.10 29.24 37.04 28.74
ENVIDR 31.22 22.99 29.55 32.17 29.52 21.57 31.44 26.60
Plenoxels 33.98 25.35 34.10 33.26 29.14 29.62 36.43 31.83
Gaussian Splatting 35.82 26.17 35.69 35.34 30.00 30.87 37.67 34.83
Editable
GaMeS (our) 35.38 26.03 35.89 37.16 29.62 31.55 37.56 35.12
GaMeS (our - pseudo-mesh) 34.74 25.94 35.88 36.58 30.50 30.66 37.84 35.41
RIP-NeRF 34.84 24.89 33.41 34.19 28.31 30.65 35.96 32.23
SSIM $\uparrow$
Static
NeRF 0.967 0.925 0.961 0.980 0.949 0.856 0.974 0.964
VolSDF 0.949 0.893 0.951 0.969 0.954 0.842 0.972 0.929
Ref-NeRF 0.974 0.929 0.975 0.983 0.921 0.864 0.979 0.954
ENVIDR 0.976 0.930 0.961 0.984 0.968 0.855 0.963 0.987
Plenoxels 0.977 0.933 0.975 0.985 0.949 0.890 0.980 0.976
Gaussian Splatting 0.987 0.954 0.983 0.991 0.960 0.907 0.985 0.987
Editable
GaMeS (our) 0.987 0.953 0.982 0.992 0.952 0.904 0.985 0.986
GaMeS (our - pseudo-mesh) 0.983 0.945 0.981 0.992 0.958 0.888 0.985 0.986
RIP-NeRF 0.980 0.929 0.977 0.962 0.943 0.916 0.963 0.979
LPIPS $\downarrow$
NeRF 0.046 0.091 0.050 0.028 0.063 0.206 0.121 0.044
VolSDF 0.056 0.119 0.054 0.191 0.048 0.191 0.043 0.068
Ref-NeRF 0.029 0.073 0.025 0.018 0.078 0.158 0.028 0.056
Plenoxels 0.031 0.067 0.028 0.015 0.057 0.134 0.037 0.026
ENVIDR 0.031 0.080 0.054 0.021 0.045 0.228 0.072 0.010
Gaussian Splatting 0.012 0.037 0.016 0.006 0.034 0.106 0.020 0.012
Editable
GaMeS (our) 0.009 0.038 0.014 0.005 0.042 0.090 0.017 0.012
GaMeS (our - pseudo-mesh) 0.012 0.039 0.016 0.006 0.036 0.109 0.018 0.012
RIP-NeRF - - - - - - - -
Table 6: Quantitative comparisons (PSNR) on a NeRF-Synthetic dataset showing that GaMeS gives comparable results with other models.

### v4

@@ 3 GaMeS : Mesh-Based Gaussian Splatting
GaMeS (our tr) 34.74 25.94 35.88 36.58 30.50 30.66 37.84 35.41
Table 1: Quantitative comparisons (PSNR) on a NeRF-Synthetic dataset showing that GaMeS gives comparable results with other models. GaMeS tr means Triangle Soup.
As the first scenario in our experiments, we choose the number $k\in\mathbb{N}$ of Gaussian components per mesh face and define how densely they should be placed on its surface. For mesh containing $n\in\mathbb{N}$ faces, the final number of Gaussians is fixed and equal to $k\cdot n$.
When we do not have mesh, we train GS, and for each Gaussian, we estimate a triangle. Then we obtain set $\mathcal{G}_{\mathcal{M}}$ of 3D parameterized Gaussians $\hat{\mathcal{G}}_{V}$ on Triangle Soup $\hat{\mathcal{M}}$:
During training, we can modify mesh vertices and Gaussian parameters. We can also use external tools for mesh fitting [21] and train only Gaussian components. In the experimental section, we show both of these scenarios.
@@
GaMeS parameterize all Gaussian components by vertices of mesh in such a way that they lie on meshes (or define Triangle Soup). Therefore, when we modify mesh/Triangle Soup, we automatically adapt the means and covariances of relevant Gaussians. Such an approach allows us to work with transformation. To better visualize such properties, we compare our model with a solution where the Gaussian components’ centers are placed on mesh faces but do not parameterize the covariance matrix. In Fig. 7, we show that GaMeS Gaussian component adapts automatically to mesh modification.
@@ 4 Experiments
Figure 9: A diagram illustrating the model, covering two scenarios: (1) with a pre-existing mesh, which may either be well-fitted and left unmodified during training (baseline) or serve as a rough initialization like FLAME model that is refined throughout training, and (2) without an initial mesh or mesh estimation provided as input.
Here, we delineate implementation details and describe the used datasets and the reasons for their selection. In our experiments, we consider two distinct scenarios, i.e., mesh-initialized input, and mesh-free input (see Fig. 9).
Implementation Details and Data Description
The time taken by the optimization process for our model varied from a few up to no more than sixty minutes, as it depends on factors like scene complexity, dataset characteristics, the number of Gaussians per face, and vertex fine-tuning decisions. Nonetheless, all experiments were conducted within a reasonable time frame, underscoring the efficiency of our approach across diverse scenarios. The supplementary material provides all the relevant details, and the source code is available in GitHub [^2] https://github.com/waczjoan/gaussian-mesh-splatting. We illustrated the fundamental gains of GaMeS model through experiments with three distinct datasets. These datasets are as follows.
@@
SuGaR 23.14 - 25.36 24.70 - 30.03 27.62 29.56 30.51
GaMeS 24.99 21.27 27.22 26.54 22.39 31.52 28.92 31.12 32.09
@@
Table 2: The quantitative comparisons of reconstruction capability (PSNR) on Mip-NeRF360 dataset. R-SuGaR-15K uses 1M vertices. GaMeS uses Triangle Soup.
@@
Synthetic Dataset: A NeRF-provided eight geometrically complex objects with realistic non-Lambertian materials [15]. In addition to containing images of the objects, the dataset also incorporates corresponding meshes. We leveraged this dataset for our initial experiments to unequivocally showcase the capabilities of our model and underscore its flexibility in mesh editing.
@@
Mip-NeRF360 Dataset: A dataset comprising five outdoor and four indoor scenes, each featuring intricate central objects or areas against detailed backgrounds [2]. We used this dataset to demonstrate the adaptability of our model in handling dense scenes, employing a non-conventional strategy during the initial training stage in mesh preparation.
@@
Human Faces Dataset: A modified subset of the Stirling/ESRC 3D Face Database [^3] https://pics.stir.ac.uk/ESRC/index.htm, that includes images of six people, generated using Blender software from various perspectives, consistently excluding the backs of heads. Only one expression is assigned for each face. Notably, the dataset lacks corresponding mesh representations for the depicted person. We used this dataset to compare GaMeS with the NeRFlame model [25], mainly within the human-mesh fitting task.
Scenario I: mesh-initialized input
Original mesh Subdivide larger faces
Refer to caption Refer to caption
Figure 10: GaMeS (baseline) uses a fixed number of faces. In the case of a mesh containing faces of different sizes, finding such parameters causes problems. In such cases, we propose to divide large faces into smaller parts. The left side includes only the original mesh, while on the right, sizable mesh faces of the brown pool surrounding have been subdivided into smaller ones.
In this scenario, we utilize the provided mesh, incorporating Gaussians by strategically placing them onto its surface. We also allowed for moving vertices. In implementation, we use a fixed number of Gaussian components per face. In the case of models with various face sizes, finding such parameters is not trivial. To solve such a problem, we propose to divide large faces into smaller parts, as shown in Fig. 10.
@@
Well-fitted pre-existing mesh (baseline): Our initial experiments used the Synthetic Dataset [15], incorporating shared meshes. Table 1 presents PSNR metrics demonstrating the competitiveness of GaMeS approach with existing methods. We provide additional numerical comparisons in the supplementary material. Notably, a significant contribution lies in the ease with which renders can be modified manually or in an automated process. Fig. 1 shows examples of reconstruction and simple object animations.
@@
Poor-fitted pre-existing mesh: GaMeS with FLAME: In this experiment, we show the possibility of fitting mesh using a pre-existing mesh, which was not acquired directly from our data in contrast to previous sub-scenarios. Here, we relied on the Faces Dataset as it provides the required experimental data. To initialize the meshes, we used the FLAME [13] framework, generating fully controllable meshes of human faces. Consequently, we could highlight the critical advantage of GaMeS model, i.e., the ability to modify objects. We used the official implementation of the FLAME, with the number of parameters suggested in RingNet [20] In addition, we implemented an extra detail learning parameter to better represent hair or T-shirts.
Initially, a human mesh is generated using the FLAME model. By utilizing the FLAME parameters, we can modify various aspects, especially facial expressions, like smiling, sadness, eye closure, mouth opening, and flirty or disgusted expressions. The possibilities of reconstruction and modifications for all six faces are depicted in Fig. 11. All expressions were chosen randomly as our main goal was to show GaMeS model’s ability to render objects as well as its modifications/animations.
Table 3 shows the results of using PSNR metrics on six faces. We made the comparisons between the two baseline NeRF models, GS, and the models created from them, i.e., the corresponding NeRFlame and GaMeS with animation capabilities.
Scenario II: mesh-free input
When we do not have mesh, we train GS with flat Gaussians. Then, we construct the Triangle Soup. In Table 1, we show a comparison with other models on the Synthetic Dataset. Our approach is also capable of modeling unbounded scenes. In order to achieve that, we first create an initial Triangle Soup and then adjust it during the training. Results in Table 2 demonstrate that such modeling is feasible and reasonably comparable to other existing models. Particularly, the results show parity with the GS-30K and R-SuGaR-15K models.
In this work, we demonstrate the editability of Triangle Soup. Fig. 1 in the first and second rows show that edits to the ficus’s or excavator’s Triangle Soup can yield highly realistic modifications, whereas the third row illustrates hotdog’s edits that are unrealistic. In Fig. 2, we explore three distinct approaches to editing a Ficus model. The first method involves direct modifications to the Triangle Soup, while the second and third methods employ deformations via an estimated mesh, as proposed in [23]. Meshes were scaled and structured using the AlphaShape algorithm [7], facilitating a comparison between realism and edit complexity. We observe that increased accuracy in mesh estimation enables more intricate edits, further expanding the range of achievable modifications.
@@ 5 Conclusion
By integrating these two concepts, we enable a structured approach to scene modeling that connects seamlessly to a pre-existing mesh. Our results demonstrate that the Triangle Soup is well-suited for replacing conventional meshes in animation contexts, allowing a flexible parametrization of a Triangle Soup based on the estimated mesh. Leveraging the capabilities of the GaMeS model, we can parametrically control flat Gaussians, facilitating intuitive and precise real-time object editing.
@@
Limitations GaMeS allows real-time modifications, but artifacts appear in case of significant changes in the case of meshes with large faces. In practice, large faces should be divided into smaller ones. How to change Gaussian components in GaMeS when mesh faces are split is not apparent.
Social Impact The model has the potential to significantly enhance object modification in games and virtual reality, offering exciting applications in simulations, particularly those involving physics-based interactions.